40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Mar 11 18:43:07 2018, Craig, DailyProgress, PEM, Notes from today's work 
    Reply  Mon Mar 12 01:25:23 2018, Craig, DailyProgress, PEM, Notes from today's work 
       Reply  Mon Mar 12 15:22:42 2018, Craig, Jaime, DailyProgress, PEM, Notes from today's work 
    Reply  Thu Mar 15 10:29:04 2018, awade, DailyProgress, scatter, Addressing 500 Hz scatter pickup 2018-03-14_16.58.21.jpg2018-03-14_20.40.18.jpg2018-03-14_16.58.13.jpg
Message ID: 2129     Entry time: Sun Mar 11 18:43:07 2018     Reply to this: 2130   2135
Author: Craig 
Type: DailyProgress 
Category: PEM 
Subject: Notes from today's work 

Today I buzzed the table and determined there was a strong 500 Hz dirty resonance on the first steering mirror after the PMC. 
This caused me to go around tightening bolts everywhere, including the offending steering mirror and the optics around it.  This did not reduce the resonance.
I tightening the PMC REFL steering mirror as well, and this caused a misalignment onto the PMC REFL PD.  This took me a little while to figure out why the North path refused to lock.  I realigned the PMC REFL steering mirror into the PD.
After I got the North PMC locking again, the North path itself was not locking anymore.  I reranged the autolocker slow volts, but this did not help. 
Turns out the North Trans PD threshold voltage was not high enough.  This is likely because of the bolt tightening, causing some slight misalignment into the North cavity, lowering the overall circulating power in the cavity.  I lowered the autolocker threshold from 1.1 volts to 1.0 volts, and aligned the North Trans PD.  We need to rescan the North cavity to get better alignment/mode matching, but I'm gonna put this off until we replace this offending 500 Hz post-PMC steering mirror.
While I was realigning the Trans PD, I noticed that even touching the trans optics tables causes large ~1Hz oscillations in the trans voltage.  This is definitely exacerbating any scattering problem we have.  Also, the Trans PD output for both paths is "breathing", going up and down with a period of about a minute.  This is bad for our autolocker's threshold.  It's possible that we should build two periscopes for the north and south paths to eliminate these elevated tables which cause coherent oscillations on all trans optics.  We could copy Tara's front periscope design.

ELOG V3.1.3-